OpenStudy (anonymous):

A chain store manager has been told by the main office that daily profit, P, is related to the number of clerks working that day, x, according to the equation P = -25x2 + 450x. What number of clerks will maximize the profit?

OpenStudy (lxelle):

P(x) = -25x^2 + 450x differentiate - P'(x) = -50x + 450

OpenStudy (igreen):

dP/dx = -50x + 300 To find the max (or min), is where the gradient (dp/dx) is zero, so ask that... 0 = -50x + 300 hence x = 6. Prove it is a max too, by differentiating again...d2P/dx2 = -50. This is -ve therefore MAX So max number of clerks = 6. This maximises the profit, P to be \(-25*(6)^2+300*6\) max profit = 900 http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.2375.html

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @desirae_michelle17

OpenStudy (anonymous):

What number of clerks will maximize the profit?

OpenStudy (anonymous):

6 is not right

OpenStudy (lxelle):

9

OpenStudy (igreen):

Another source says 9.

OpenStudy (igreen):

Lol.

OpenStudy (anonymous):

thank you so much , 9 is correct

OpenStudy (lxelle):

:)

OpenStudy (lxelle):

do you know how to get 9?

OpenStudy (anonymous):

yesss i solved it out . i had to show work

OpenStudy (lxelle):

great.